# A Survey for Deep Reinforcement Learning Based Network Intrusion Detection

### v1

@@ VI Conclusion
Network intrusion detection has attracted much attention in broader cybersecurity in recent years. This paper presents a brief survey focused on the application of DRL techniques in the domain of network intrusion detection over the past few years. Current DRL-based network intrusion detection systems is still facing many challenges, including long time for model training, low training efficiency, and a lack of real-world deployment and evaluation. Furthermore, DRL-based NID seems to lag behind the developments in the mainstream DRL community, yet it holds significant potential for growth. Many novel DRL algorithms have not been extended to the application in NID. Additionally, we advocate for the use of data from real-world attack scenarios to train DRL-based NID systems and call for preliminary explorations into modelling attackers’ behaviours using inverse reinforcement learning and also advocate for researchers focusing on the intrusion detection for internet of things by collecting more suitable datasets. Lastly, we discussed how the world models built on large language models, or even broader large models can be used to train defensive agents, further facilitating the development of generative-augmented DRL for advancing DRL in the field of NID.

### v2

@@ VI Conclusion
Network intrusion detection has attracted much attention in broader cybersecurity in recent years. This paper presents a brief survey focused on the application of DRL techniques in the domain of network intrusion detection over the past few years. Current DRL-based network intrusion detection systems are still facing many challenges, including long time for model training, low training efficiency, and a lack of real-world deployment and evaluation. Furthermore, DRL-based NID seems to lag behind the developments in the mainstream DRL community, yet it holds significant potential for growth. Many novel DRL algorithms have not been extended to the application in NID. Additionally, we advocate for the use of data from real-world attack scenarios to train DRL-based NID systems and call for preliminary explorations into modelling attackers’ behaviours using inverse reinforcement learning and also advocate for researchers focusing on the intrusion detection for internet of things by collecting more suitable datasets. Lastly, we discussed how the world models built on large language models, or even broader large models can be used to train defensive agents, further facilitating the development of generative-augmented DRL for advancing DRL in the field of NID.
Ethical Statement
This research is a survey study that reviews and analyzes existing published literature. It does not involve any human participants, animal subjects, or the collection of personal data. Therefore, no ethical approval was required for this work.
Data Availability Statement
This article is a survey paper and does not generate any new datasets. All datasets discussed in this work are publicly available benchmark datasets, and their sources and access links have been cited and referenced within the manuscript accordingly.